OpenStudy (anonymous):

? quantity x squared plus 2 x over quantity 12 x plus 54 end quantity minus quantity 3 minus x over quantity 8 x plus 36 & 2 over quantity x squared plus 8 x plus 15 end quantity plus 1 over quantity x squared plus 11 x plus 30

OpenStudy (anonymous):

2) \[\frac{ 2 }{ x^2\div18x +15 }+\frac{ 1 }{ x^2+11x+30 }\]

OpenStudy (anonymous):

1) \[\frac{ x^2+2x }{ 12x+54 }-\frac{ 3-x }{ 8x+36 }\]